# Anti-rotors and algebra invariants

### v36

@@ 4.2. Special unital norms
The requisite open ball in the theorem statement always exists, as was already observed immediately following the statement of Definition 3.1. Let $\eta$ be such a neighborhood of $\mathbf{1}$.
Given $\alpha>0$ such that $\frac{1}{\alpha}s$ and $\alpha s$ are both in $\eta$, (4.2) implies,
@@
Let $\eta^{-1}$ be the set consisting of the multiplicative inverses of all members of $\eta$. It is easily shown that $\eta^{-1}$ is an open neighborhood of $\mathbf{1}$, and $\eta$, $\eta^{-1}$ are diffeomorphic via the multiplicative inversion operation. We now consider a smooth path $\mathcal{P}$ from $\mathbf{1}$ to $s$ which is inside $\eta$. Let $\mathcal{P}^{-1}$ be the point set consisting of the multiplicative inverses of the members of $\mathcal{P}$. Clearly, $\mathcal{P}^{-1}$ is a smooth path contained in $\eta^{-1}$. We then have,
@@
As also formalized in Section 4.4, this program additionally generalizes the “usual” algebra norm (the determinant of an element’s image under the left regular representation, e.g., [8]), since the latter is closely related to one of the $\ell_{\rm sp}(s)$. Keeping in mind details of the proof of Proposition 3.4, this can begin to be appreciated by observing that if the algebra in question is $M_{n}(\mathbb{R})$ (for which $\|\mathbf{1}_{M_{n}(\mathbb{R})}\|^{2}=n$), and $L$ is taken to be the matrix transpose operation $\mathcal{T}$ (which is a normalized uncurling metric as above), then (3.5) indicates that the corresponding solution to (4.1) is given by $\ell_{\rm sp}(s)=(\text{det}(s))^{\frac{1}{n}}$.
@@ 4.3. The geometries associated with an algebra
According to Proposition 4.3, $\ell_{\rm sp}(s^{-1})=\left(\ell_{\rm sp}(s)\right)^{-1}=\frac{1}{\ell_{\rm sp}(s)}$, and $\ell_{\rm sp}(s)$ is a degree-1 positive homogeneous function. Using both those properties, and assuming $L_{\shortparallel}$ is nonsingular, (4.1) then implies
@@
The above equation set has an obvious similarity to an equation set that holds for a quadratic space. Thus, if $L$ is the scalar product matrix associated with a quadratic space where the associated quadratic form is positive-definite or indefinite but nondegenerate, it follows that
@@
This equation indicates a reason for the name “quadratic space”. Comparing (4.13) with the first equality in (4.1) justifies referring to the algebra-associated space described by (4.9), (4.10) as a “logarithmic space”.
@@
(4.14) $\big(\ell_{\rm qs}(s)\big)^{2}=2\int_{0}^{s}(L_{\rm sym}t)\cdot\mathbf{d}t=2\int_{0}^{s}t^{T}L_{\rm sym}\mathbf{d}t.$
@@
However, any uncurling metric $L$ results in “lengths” and “angles” in the same way as the latter are obtained with $L_{\shortparallel}$. We will now show that the above implied collection of geometries associated with an algebra persists under an algebra isomorphism. That is, the respective sets of element decompositions (4.9), (4.10) only differ by a linear change of variables.
@@
If $A_{1}$ and $A_{2}$ are isomorphic algebras, then $\mathfrak{u}_{A_{1}}^{d}$ and $\mathfrak{u}_{A_{2}}^{d}$ are the same up to a change of basis.
@@ 4.4. Generalization of the usual algebra norm of an element
The usual algebra norm is given by the determinant of an element’s image under the left regular representation [8]. The dual of the anti-rotor (as the group of unital norms) subsumes this definition in the sense of Theorem 4.7, as derived in this subsection.
@@
Hence, there is at most one possible normalized uncurling metric. But according to observations made in the first paragraph following the proof of Proposition 3.4, $M_{n}(\mathbb{R})$ does indeed have a normalized uncurling metric - this being the $(n^{2}\times n^{2})$-matrix corresponding to the matrix transpose operation on $(n\times n)$-matrices. In the third paragraph following the proof of Proposition 4.3 it was shown that $(\mbox{det}(s))^{\frac{1}{n}}$ is the associated special unital norm. Since the usual norm of an element $s\in M_{n}(\mathbb{R})$ is $(\mbox{det}(s))^{n}$, the theorem follows. ∎
@@
The fraction $\frac{\|\mathbf{1}_{\bar{A}}\|^{2}}{n}$ is employed in (4.19) in order that $\langle\mathbf{1}_{\bar{A}},\mathbf{1}_{\bar{A}}\rangle_{\bar{A}}=\|\mathbf{1}_{\bar{A}}\|^{2}$, given that $\mathbf{1}_{M_{n}(\mathbb{R})}^{T}\mathcal{T}\,\mathbf{1}_{M_{n}(\mathbb{R})}=n=\|\mathbf{1}_{M_{n}(\mathbb{R})}\|^{2}$. Due to this isomorphism of $A$ and $\bar{A}$ as quadratic spaces (as established in (4.19)), the previously demonstrated $d((\mathcal{T}s^{-1})\cdot\mathbf{d}s)=0$ in some neighborhood of $\mathbf{1}_{A}$ in $A$ implies $d((\bar{\mathcal{T}}(\bar{s})^{-1})\cdot d\bar{s})=0$. Thus, $\bar{\mathcal{T}}$ is an uncurling metric for $\bar{A}$.
@@
(4.21) $(\bar{s})^{T}\bar{\mathcal{T}}(\bar{s})^{-1}=\frac{\|\mathbf{1}_{\bar{A}}\|^{2}}{n}(s^{T}\mathcal{T}s^{-1})=\|\mathbf{1}_{\bar{A}}\|^{2}.$
@@
(4.22) $\log\ell_{\rm sp}(\bar{s})=\frac{1}{\|\mathbf{1}_{\bar{A}}\|^{2}}\int_{\mathbf{1}_{\bar{A}}}^{\bar{s}}(\bar{\mathcal{T}}(\bar{s})^{-1})\cdot\mathbf{d}\bar{s}$ $=$ $\frac{1}{\|\mathbf{1}_{\bar{A}}\|^{2}}\int_{\mathbf{1}_{A}}^{s}\frac{\|\mathbf{1}_{\bar{A}}\|^{2}}{n}(\mathcal{T}s^{-1})\cdot\mathbf{d}s$
@@
since (4.4) is satisfied because $\Omega\bar{s}=\underline{s}$ and $\Omega(\bar{s})^{-1}=\underline{s}^{-1}$ (because $\Omega$ is an epimorphism of rings).
According to (4.2), $\bar{L}_{\shortparallel}$ is then associated with the special unital norm on $\bar{A}$ given by $\ell_{\rm sp}(\bar{s})=[\text{det}(\underline{\mathcal{I}}\Omega\bar{s})]^{\frac{1}{m}}$ (we are abusing notation here, since $\ell_{\rm sp}(\bar{s})$ is not the same as $\ell_{\rm sp}(\underline{s})$ on the left-hand-side of (4.23), as these special unital norms are associated with different algebras). That is, in evaluation of the integral resulting from (4.2), a path of integration from $\mathbf{1}_{\bar{A}}$ to $\bar{s}$ can be replaced by a path of integration from $\mathbf{1}_{\underline{A}}$ to $\underline{s}$ because of the projection implied by the integrand resulting from (4.24). Furthermore, on this path,
(4.25) $[\bar{L}_{\shortparallel}(\bar{t})^{-1}]\,\cdot\,\mathbf{d}\bar{t}$ $=$ $[(\bar{t})^{-1}]^{T}\bar{L}_{\shortparallel}\,\mathbf{d}\bar{t}=\frac{\|\mathbf{1}_{\bar{A}}\|^{2}}{\|\mathbf{1}_{\underline{A}}\|^{2}}[(\bar{t})^{-1}]^{T}\Omega^{T}\underline{L}_{\shortparallel}\,\Omega\,\mathbf{d}\bar{t}=\frac{\|\mathbf{1}_{\bar{A}}\|^{2}}{\|\mathbf{1}_{\underline{A}}\|^{2}}\,[\underline{t}^{-1}]^{T}\,\underline{L}_{\shortparallel}\,\mathbf{d}\underline{t}$
@@ 5. Combinatorial issues
    For which algebras is ${\rm det}(M_{\mathfrak{u}_{A}})$ identically 0?
@@ 8. How large is the class of invariants implied by this approach?
While we have derived a mechanism that generates algebra isomorphism invariants based on selection of $f(s)=s^{-1}$ in Definition 3.1, there are infinitely many relevant possibilities for $f(s)$ that could substitute for $s^{-1}$. One might then wonder if the association of an algebra with its resulting anti-rotor “variants” (each resulting from a different choice of $f(s)$) could thereby be an endless source of algebra invariants. If so, it would be appropriate to wonder how one could bring order to a such a collection of processes that literally spew out invariants. Would there be different classes of functions that would generate different classes of invariants? And so on.
@@
But before contemplating this issue too deeply, we observe that there is one thing more we can glean from the fourteen rows of algebras in Table III. If in Definition 3.1 we replace $s^{-1}$ with $s^{j}$ for small $|j|$, $j\in\mathbb{Z}$, $j\neq 0,1$ (i.e., excluding the cases that are not dependent on the algebra product), it is quickly discovered that all resulting anti-rotor variants for the algebras in Table III are identical within each choice of algebra. Extrapolating, if the above in fact holds for all $j\in\mathbb{Z}$, $j\neq 0,1$, linearity of the exterior derivative means it would hold for arbitrary linear combinations of the various $s^{j}$ - implying that with respect to suitable domains in $\mathbb{R}^{n}$ it would hold for functions expressible by infinite series in the powers of $s$.
@@
However, the behavior cited in the first sentence of the prior paragraph was unanticipated and seems very surprising. There are several obvious avenues one might think would lead to an understanding of it.
First, let $\hat{s}$ be the left regular representation of an algebra with elements $s$, so that we can write $s^{j}=\hat{s}^{j-1}[s]$, where $s$ ranges over the algebra elements in an appropriate neighborhood of $\mathbf{1}$. For any uncurling metric $L$ (i.e., such that $Ls^{-1}$ is irrotational), extrapolation of the behavior described in the penultimate paragraph suggests that the various vector fields defined by $Ls^{j}=(L\hat{s}L^{-1})^{j-1}Ls$ are simultaneously irrotational. That is, repetitive applications of $L\hat{s}L^{-1}$ or $L\hat{s}^{-1}L^{-1}$ initially seeded with the irrotational vector field $Ls^{-1}$ always result in irrotational vector fields.
@@
One might incorrectly suspect that this follows because repeated applications of an operator that resulted in an irrotational vector field will continue to result in irrotational vector fields. However, consider the vector field $(F(s))[s]$ where $F(s)$ is an $(n\times n)$-matrix whose entries are smooth functions of components of $s$, and suppose $(F(s))[s]$ is irrotational. In general, this does not imply that $(F(s))^{j}[s]$ is irrotational for all $j\in\mathbb{Z}$. For example, let $s=\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$, and $F(x,y)=\left(\begin{array}[]{ccc}x&2y\\ 2y&2x\end{array}\right)$. Then it is easily checked that $(F(x,y))\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$ is irrotational and, in fact, $(F(x,y))^{2}\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$ is also irrotational. But $(F(x,y))^{3}\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$ is not irrotational. Of course, an important distinction between $L\hat{s}L^{-1}$ and $F(x,y)$ is that the former are elements of an algebra and the latter are not. So, generalization of the phenomenon described in the second paragraph of this section, if valid, would seem to depend on the presence of an algebra. But how would that work?
@@
Another possibility for understanding the behavior cited in the second paragraph of this subsection is suggested by what happens in the case of algebra $M_{n}(\mathbb{R})$, with respect to which there is the well known Matrix Calculus expression,
(8.1) $\frac{\partial}{\partial s}\left(\text{Tr}(s^{j})\right)=j\left(s^{j-1}\right)^{T},$
@@
where Tr indicates trace (in the Matrix Calculus literature, $s$ in (8.1) is typically written as $\mathbf{X}$). Using Jacobi’s formula, we have noted earlier that the transpose operation T is an uncurling metric for $M_{n}(\mathbb{R})$ in the exceptional $f(s)=s^{-1}$ case, and (8.1) indicates that T is indeed also an uncurling metric for $f(s)=s^{j}$ in general. So, according to the proof of Proposition 4.5, here we have an example of an algebra whose space of uncurling metrics (in this case, one-dimensional) is independent of the choice of $f(s)=s^{j}$, for $j\neq 0,1$. Now, if we deal with the left regular representation of an algebra, then we will be working with subalgebras of $M_{n}(\mathbb{R})$ (and their usually larger spaces of uncurling metrics). In the latter setting, perhaps there is there an analogy to be exploited as regards the suggestive (8.1). But how would that work?
@@
As a third way to try to find an explanation for the behavior cited in the second paragraph of this subsection, suppose it can be shown that a pullback deriving from a diffeomorphism $\phi_{\mu,j}$ exists between the 1-forms $(\mu s^{2})\cdot\mathbf{d}s$ and $(\mu s^{j})\cdot\mathbf{d}s$ when either of them is exact - i.e., whenever $\mu$ is an uncurling metric for either $s^{2}$ or $s^{j}$, $j\neq 0,1$. That is, we would have $(\mu s^{j})\cdot\mathbf{d}s=\phi_{\mu,j}^{*}\big((\mu s^{2})\cdot\mathbf{d}s\big)$, along with a similar equation expressing $\big(\mu s^{2}\big)\cdot\mathbf{d}s$ as a pullback of $\big(\mu s^{j}\big)\cdot\mathbf{d}s$. Commutativity of a pullback and a differential form would then imply,
$d\left((\mu s^{j})\cdot\mathbf{d}s\right)=d\left(\phi_{\mu,j}^{*}(\mu s^{2})\cdot\mathbf{d}s\right)=\phi_{\mu,j}^{*}d\big((\mu s^{2})\cdot\mathbf{d}s\big).$
This would mean that the exterior derivative of $(\mu s^{j})\cdot\mathbf{d}s$ vanishes if and only if the exterior derivative of $(\mu s^{2})\cdot\mathbf{d}s$ vanishes - and we would have our explanation.
@@
However, proving that those pullbacks exist would involve a demonstration of the existence of solutions to large algebra-dependent collections of nonlinear partial differential equations. For individual cases of a particular algebra, a particular uncurling metric $\mu^{\prime}$, and a particular value of $j$, along with the (unwelcome) stipulation that the diffeomorphisms must have only rational function components, the problem might be addressed with known techniques of computational algebraic geometry. But it is difficult to overcome the impression that the general problem is intractable, i.e., that of establishing the existence of the requisite pullbacks for any algebra and any $j\neq 0,1$.
Thus, the problem of explaining the behavior cited in the second paragraph of this subsection appears daunting. So we are left with the following, which for convenience addresses the left regular representations of algebras.

### v37

@@ 4.2. Special unital norms
The requisite open ball in the theorem statement always exists, as was already observed immediately following the statement of Definition 3.1. Let domain $D$ be such a neighborhood of $\mathbf{1}$.
Given $\alpha>0$ such that $\frac{1}{\alpha}s$ and $\alpha s$ are both in $D$, (4.2) implies,
@@
Define $D^{-1}$ to be the set consisting of the multiplicative inverses of all members of $D$. It is easily shown that $D^{-1}$ is an open neighborhood of $\mathbf{1}$, and the domains $D$, $D^{-1}$, are diffeomorphic via the multiplicative inversion operation. We now consider a smooth path $\mathcal{P}$ from $\mathbf{1}$ to $s$ which is inside $D$. Let $\mathcal{P}^{-1}$ be the point set consisting of the multiplicative inverses of the members of $\mathcal{P}$. Clearly, $\mathcal{P}^{-1}$ is a smooth path contained in $D^{-1}$. We then have,
@@
As also formalized in Section 4.4, this program additionally generalizes the “usual” algebra norm (the determinant of an element’s left regular representation, e.g., [8]), since the latter is closely related to one of the $\ell_{\rm sp}(s)$. Keeping in mind details of the proof of Proposition 3.4, this can begin to be appreciated by observing that if the algebra in question is $M_{n}(\mathbb{R})$ (for which $\|\mathbf{1}_{M_{n}(\mathbb{R})}\|^{2}=n$), and $L$ is taken to be the matrix transpose operation $\mathcal{T}$ (which is a normalized uncurling metric as above), then (3.5) indicates that the corresponding solution to (4.1) is given by $\ell_{\rm sp}(s)=(\text{det}(s))^{\frac{1}{n}}$.
@@ 4.3. The geometries associated with an algebra
As discussed in [6], George Birkhoff showed that his four-postulate formulation of Euclidean geometry implies that the plane is an affine space whose vector space component is the Cartesian plane [9]. Despite the fact that his formulation is in principle analytic (since it intrinsically incorporates $\mathbb{R}$), as preparation he provided what is essentially a synthetic proof of the Pythagorean Theorem. Ultimately, according to his program, the theorem is equivalent to the statement that $\mathbb{R}\oplus\mathbb{R}$ is endowed with a length function satisfying $\ell(x,y)=\sqrt{x^{2}+y^{2}}$. As observed in [6], for $s=(x,y)\neq(0,0)$, it is a simple matter to demonstrate that the equation in the prior sentence is itself equivalent to,
$s=\ell(s)\nabla\ell(s),$
$\ell(\nabla\ell(s))=1.$
So we could instead take the above characterization of the non-origin points of $\mathbb{R}\oplus\mathbb{R}$ to be the Pythagorean Theorem. If one considers the format of this characterization to be the important feature, we might deem the following (and the suggested appellation) as faithful to its essence.
Corollary 4.4 (The Pythagorean Theorem for algebras).
For a given algebra, suppose $L_{\shortparallel}$ is a nonsingular normalized uncurling metric and $\ell_{\rm sp}(s)$ is its dual. Then for elements $s$ in a unit ball around $\mathbf{1}$ devoid of non-units,
@@
Proof.
According to Proposition 4.3, $\frac{1}{\ell_{\rm sp}(s)}=\ell_{\rm sp}(s^{-1})$. Applying this to the right-hand-side of (4.1), next applying $L_{\shortparallel}^{-1}$ to both sides of the resulting equation, and then substituting $s^{-1}$ for $s$, we obtain (4.9). Now applying $\ell_{\rm sp}$ to both sides of the latter equation, and invoking the degree-1 positive homogeneity of $\ell_{\rm sp}$ that is another consequence of Proposition 4.3, we obtain (4.10). ∎
The equation set in Corollary 4.4 has an even more obvious similarity to an equation set that holds not just for Euclidean space, but for quadratic spaces in general. Thus, if $L$ is the scalar product matrix associated with a quadratic space where the associated quadratic form is positive-definite or indefinite but nondegenerate, it follows that
@@
The exponent on the right-hand-side indicates the reason for the name “quadratic space”. Comparison of (4.13) with the first equality in (4.1) justifies referring to the algebra-associated space described by (4.9), (4.10) as a “logarithmic space”.
@@
(4.14) $\ell_{\rm qs}^{2}(s)=2\int_{0}^{s}(L_{\rm sym}t)\cdot\mathbf{d}t=2\int_{0}^{s}t^{T}L_{\rm sym}\mathbf{d}t.$
@@
However, any uncurling metric $L$ results in “norms” and “angles” in the same way as the latter are obtained with $L_{\shortparallel}$. The following shows that the implied collection of geometries associated with an algebra $A$ (i.e., resulting from the members of $\mathfrak{u}_{A}^{d}$ as above) persists under an algebra isomorphism.
@@
If $A_{1}$ and $A_{2}$ are isomorphic algebras, then the respective collections of geometries implied by $\mathfrak{u}_{A_{1}}^{d}$ and $\mathfrak{u}_{A_{2}}^{d}$ are the same.
@@
Obviously, the theorem does require the isomorphism assumption. An arbitrary basis change applied to $A_{1}$ would not in general result in the dual of the anti-rotor of some other algebra $A_{3}$ such that $A_{3}$ is isomorphic to $A_{1}$.
@@ 4.4. Generalization of the usual algebra norm of an element
As already noted, the usual algebra norm is given by the determinant of an element’s left regular representation. The dual of the anti-rotor (as the group of unital norms) subsumes this definition in the sense of Theorem 4.8, as derived in this subsection.
@@
Hence, there is at most one possible normalized uncurling metric. But according to observations made in the penultimate paragraph of Section 4.2, $M_{n}(\mathbb{R})$ does indeed have a normalized uncurling metric - this being the $(n^{2}\times n^{2})$-matrix $\mathcal{T}$ corresponding to the matrix transpose operation on $(n\times n)$-matrices. In the final paragraph of Section 4.2 it was shown that $(\mbox{det}(s))^{\frac{1}{n}}$ is the associated special unital norm. Since the usual norm of an element $s\in M_{n}(\mathbb{R})$ is $(\mbox{det}(s))^{n}$, the theorem follows. ∎
@@
The fraction $\frac{\|\mathbf{1}_{\bar{A}}\|^{2}}{n}$ is employed in (4.19) in order that $\langle\mathbf{1}_{\bar{A}},\mathbf{1}_{\bar{A}}\rangle_{\bar{A}}=\|\mathbf{1}_{\bar{A}}\|^{2}$, given that $\mathbf{1}_{M_{n}(\mathbb{R})}^{T}\mathcal{T}\,\mathbf{1}_{M_{n}(\mathbb{R})}=n=\|\mathbf{1}_{M_{n}(\mathbb{R})}\|^{2}$. Due to this isomorphism of $A$ and $\bar{A}$ as quadratic spaces (as established in (4.19)), the previously demonstrated $d((\mathcal{T}s^{-1})\cdot\mathbf{d}s)=0$ in some neighborhood of $\mathbf{1}_{A}$ in $A$ implies $d((\bar{\mathcal{T}}\bar{s}^{-1})\cdot d\bar{s})=0$. Thus, $\bar{\mathcal{T}}$ is an uncurling metric for $\bar{A}$.
@@
(4.21) $\bar{s}^{T}\bar{\mathcal{T}}\bar{s}^{-1}=\frac{\|\mathbf{1}_{\bar{A}}\|^{2}}{n}(s^{T}\mathcal{T}s^{-1})=\|\mathbf{1}_{\bar{A}}\|^{2}.$
@@
(4.22) $\log\ell_{\rm sp}(\bar{s})=\frac{1}{\|\mathbf{1}_{\bar{A}}\|^{2}}\int_{\mathbf{1}_{\bar{A}}}^{\bar{s}}(\bar{\mathcal{T}}\bar{s}^{-1})\cdot\mathbf{d}\bar{s}$ $=$ $\frac{1}{\|\mathbf{1}_{\bar{A}}\|^{2}}\int_{\mathbf{1}_{A}}^{s}\frac{\|\mathbf{1}_{\bar{A}}\|^{2}}{n}(\mathcal{T}s^{-1})\cdot\mathbf{d}s$
@@
since (4.4) is satisfied because $\Omega\bar{s}=\underline{s}$ and $\Omega\bar{s}^{-1}=\underline{s}^{-1}$ (because $\Omega$ is an epimorphism of rings).
According to (4.2), $\bar{L}_{\shortparallel}$ is then associated with the special unital norm on $\bar{A}$ given by $\ell_{\rm sp}(\bar{s})=[\text{det}(\underline{\mathcal{I}}\Omega\bar{s})]^{\frac{1}{m}}$ (we are abusing notation here, since $\ell_{\rm sp}(\bar{s})$ is obviously not the same function as $\ell_{\rm sp}(\underline{s})$ on the left-hand-side of (4.23), as these special unital norms are associated with different algebras - but it is convenient to use the same notation for them, and no confusion arises). That is, in evaluation of the integral resulting from (4.2), a path of integration from $\mathbf{1}_{\bar{A}}$ to $\bar{s}$ can be replaced by a path of integration from $\mathbf{1}_{\underline{A}}$ to $\underline{s}$ because of the projection implied by the integrand resulting from (4.24). Furthermore, on this path,
(4.25) $[\bar{L}_{\shortparallel}\,\bar{t}^{-1}]\,\cdot\,\mathbf{d}\bar{t}$ $=$ $\left(\bar{t}^{-1}\right)^{T}\bar{L}_{\shortparallel}\,\mathbf{d}\bar{t}=\frac{\|\mathbf{1}_{\bar{A}}\|^{2}}{\|\mathbf{1}_{\underline{A}}\|^{2}}\left(\bar{t}^{-1}\right)^{T}\Omega^{T}\underline{L}_{\shortparallel}\,\Omega\,\mathbf{d}\bar{t}=\frac{\|\mathbf{1}_{\bar{A}}\|^{2}}{\|\mathbf{1}_{\underline{A}}\|^{2}}\,\left(\underline{t}^{-1}\right)^{T}\,\underline{L}_{\shortparallel}\,\mathbf{d}\underline{t}$
@@ 5. Combinatorial issues
    How many algebra isomorphism classes are there such that ${\rm det}(M_{\mathfrak{u}_{A}})$ is identically 0?
@@ 8. How large is the class of invariants implied by this approach?
8.1. Type I anti-rotors
While we have derived a mechanism that generates algebra isomorphism invariants based on the specification of $f(s)=s^{-1}$ in Definition 3.1, there are infinitely many relevant possibilities for $f(s)$ dependent on the algebra’s bilinear product that could substitute for $s^{-1}$.
Definition 8.1.
For $j\in\mathbb{Z}$, $j\neq 0,1$, if $s^{j}$ substitutes for $s^{-1}$ in Definition 3.1, the set of metrics satisfying the accordingly modified (3.1) is the Type I anti-rotor associated with $f(s)=s^{j}$.
A trivial modification of the proof of Theorem 3.2 results in,
Corollary 8.2.
Given two isomorphic algebras, their respective Type I anti-rotors associated with $f(s)=s^{j}$ are the same up to a change of basis.
@@ 8.1. Type I anti-rotors
One might then wonder if the association of an algebra with the above multiplicity of anti-rotors (including those based on linear combinations of the above $s^{j}$) could thereby be an endless source of algebra invariants. If so, it would be appropriate to wonder how one could bring order to such a collection of processes that literally spew out invariants.
@@
But before contemplating this issue too deeply, we observe that there is one thing more we can glean from the fourteen rows of algebras in Table III. For each algebra, all of its Type I anti-rotors associated with the various $f(s)=s^{j}$ are identical for “small” $|j|$. Extrapolating, for any given algebra, if the above holds for all $j\in\mathbb{Z}$ except for $j=0$ or $j=1$ (by definition), linearity of the exterior derivative means it would hold for arbitrary linear combinations of the various $s^{j}$ - implying that with respect to suitable domains in $\mathbb{R}^{n}$ it would hold for functions expressible by infinite series in these powers of $s$. That is, the kernel of $\not{\text{S}}_{f}$ would be independent of $f$.
@@
However, the behavior cited in the second sentence of the prior paragraph was unanticipated and seems surprising. There are several obvious avenues one might initially think would lead to an understanding of it.
@@
First, let $\hat{s}$ be the left regular representation of an algebra with elements $s$, so that we can write $s^{j}=\hat{s}^{j-1}[s]$, where $s$ ranges over the algebra elements in an appropriate neighborhood of $\mathbf{1}$. The behavior noted in the penultimate paragraph suggests that since $Ls$ is irrotational, and assuming $(L\hat{s}L^{-1})Ls=Ls^{2}$ is irrotational, then repeated applications of $L\hat{s}L^{-1}$ will continue to produce irrotational vector fields, i.e., $Ls^{j}=(L\hat{s}L^{-1})^{j-1}Ls=Ls^{j}$ is irrotational. Why? One might suspect that this follows because, in general, when an operator applied to an irrotational vector field results in an irrotational vector field, then it will continue to produce irrotational vector fields on all successive applications - which is false. That is, consider the vector field $(F(s))[s]$ where $F(s)$ is an $(n\times n)$-matrix whose entries are smooth functions of components of $s$, and suppose $(F(s))[s]$ is irrotational. This does not imply that $(F(s))^{j}[s]$ is irrotational for all $j\in\mathbb{Z}$. For example, let $s=\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$, and $F(x,y)=\left(\begin{array}[]{ccc}x&2y\\ 2y&2x\end{array}\right)$. Then it is easily checked that $(F(x,y))\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$ is irrotational and, in fact, $(F(x,y))^{2}\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$ is also irrotational. But $(F(x,y))^{3}\left(\begin{array}[]{ccc}x\\ y\end{array}\right)$ is not irrotational.
@@
Another possibility for understanding the behavior cited in the second paragraph following Corollary 8.2, is inspired by the easily appreciated truth of the generalization of that behavior to all $j\in\mathbb{Z}$ in the particular case of the full matrix algebra $M_{n}(\mathbb{R})$. Specifically, there is the well known Matrix Calculus expression valid for all $j\in\mathbb{Z}$ and for $s$ in a neighborhood of $\mathbf{1}$,
(8.1) $\frac{\partial}{\partial s}\left(\text{Tr}(s^{j+1})\right)=(j+1)\left(s^{j}\right)^{T},$
@@
where Tr indicates trace (in the Matrix Calculus literature, $s$ in (8.1) is typically written as $\mathbf{X}$). Using Jacobi’s formula, we have noted earlier that the transpose operation T (i.e., $\mathcal{T}$) is an uncurling metric for $M_{n}(\mathbb{R})$ in the exceptional $f(s)=s^{-1}$ case for which (8.1) gives no useful information. Thus, for all other $j\in\mathbb{Z}$, (8.1) indicates that T is indeed also a member (uncurling metric) of each of the Type I anti-rotors defined by the different $f(s)=s^{j}$ (since the expression on the left-hand-side of (8.1) is a gradient). So, according to an easy extrapolation of the argument in the proof of Proposition 4.6 (extending the argument from $f(s)=s^{-1}$ to $f(s)=s^{j}$, $j\neq 0,1$, to again indicate that the only uncurling metrics for $M_{n}(\mathbb{R})$ are multiples of $\mathcal{T}$), here we have an example of an algebra whose space of uncurling metrics is indeed independent of the choice of $f(s)=s^{j}$, for $j\neq 0,1$. On the other hand, it is not obvious how this observation can be adapted to the case of subalgebras of $M_{n}(\mathbb{R})$.
As an aside, the observations in the last paragraph lead to,
Proposition 8.3.
For any given algebra, the Type I anti-rotor associated with $f(s)=s^{j}$ for any $j\in\mathbb{Z}$, $j\neq 0,1$, has nontrivial members.
Proof.
According to the penultimate paragraph, the algebra $M_{n}(\mathbb{R})$ is such that its Type I anti-rotor associated with $f(s)=s^{j}$ for any $j\in\mathbb{Z}$ has nontrivial members - resulting from scalar multiples of the transpose operator. Identical arguments as appearing in the proofs of Lemma 4.7 and Theorem 4.8, except replacing $s^{-1}$ with $s^{j}$, then establish the proposition.∎
@@
Returning to our original quest, a third way to try to find an explanation for the behavior cited in the second paragraph following Corollary 8.2, is to exploit the fact that pullbacks commute with the exterior derivative. Thus, suppose it could be shown that for any $\neq 0,1, jj\in\mathbb{Z}$, one can find a $j$-dependent diffeomorphism $\phi$ between relevant domains in $\mathbb{R}^{n}$ resulting in a $j$-dependent pullback $\phi^{*}$ between the 1-forms $(\mu s^{-1})\cdot\mathbf{d}s$ and $(\mu s^{j})\cdot\mathbf{d}s$. Then we would have $\phi^{*}\big[(\mu s^{-1})\cdot\mathbf{d}s\big]=(\mu s^{j})\cdot\mathbf{d}s$, along with a similar equation expressing $\big(\mu s^{-1}\big)\cdot\mathbf{d}s$ as a pullback of $\big(\mu s^{j}\big)\cdot\mathbf{d}s$. Commutativity of the exterior derivative and a pullback would then imply that the exterior derivative of $(\mu s^{j})\cdot\mathbf{d}s$ vanishes if and only if the exterior derivative of $(\mu s^{-1})\cdot\mathbf{d}s$ vanishes - and we would have our explanation.
@@
However, proving that all those pullbacks exist would involve a demonstration of the existence of solutions to each set of nonlinear partial differential equations arising from the choice of any given algebra. For individual cases of a particular algebra and a particular value of $j$, along with the stipulation that the relevant diffeomorphisms must have only rational function components, the problem might be addressed with known techniques of computational algebraic geometry. But it is difficult to overcome the impression that the general problem is intractable, i.e., that of establishing the existence of the requisite pullbacks for any algebra and any $j\neq 0,1$.
Thus, at this point, our limited computational evidence only leaves us with,